# User-Aware Prefix-Tuning is a Good Learner for Personalized Image Captioning

### v1

@@ 2.1 Image Captioning
Image captioning serves as a pivotal link between Computer Vision (CV) and Natural Language Processing (NLP), necessitating the generation of contextually and grammatically appropriate textual descriptions for images. Notable progress has been achieved in this field with the advent of deep learning techniques. Earlier approaches utilized convolutional neural network [10] to extract visual features from images and recurrent neural network [17] to generate captions [16, 32]. Subsequent studies introduced attention mechanisms, leading to improved caption quality [1, 35, 37]. In addition, there has been an emergence of large-scale vision and language pre-trained models that can be applied to image captioning [34, 39, 36, 27]. However, these models often require significant computational resources and training efforts. To address this issue, Mokady et al. [18] proposed a strategy of keeping the pre-trained weights fixed and focusing on training specific network components to bridge the gap between visual and textual representations. Despite the achievements of these efforts, they are limited to generating neutral captions, which may not fully cater to the diverse needs of users in the era characterized by the rapid emergence of social media.
@@ 2.2 Personalized Vision and Language Research
Driven by the rapid proliferation of social media and the growing demand for personalized content, numerous scholars have conducted research with a primary focus on customizing the visual and linguistic domains [4, 3, 5]. Shuster et al. [26] defined 215 potential personality traits to guide caption generation. However, in real-world scenarios, users’ styles may not be adequately captured by these traits. Park et al. [20] constructed datasets specifically for personalized image captioning and employed the memory network to store the users’ active vocabularies, enabling the generation of personalized captions. Long et al. [14] proposed a model for personalized image captioning that can be applied across different domains, thereby expanding the range of potential application scenarios. Considering users’ long-term and short-term stylistic preferences, Zhang et al. [40] proposed an approach that combines these temporal dimensions. However, the implementation of this approach necessitates the collection of more detailed data. To generate captions that align with user preferences, Wang et al. [33] utilized user-provided tags as a foundational element in their methodology. Zeng et al. [38] leveraged user profiles such as age and gender to generate social media comments. Our approach eliminates the need for supplementary information, such as user profiles or posting timestamps, and only requires training a small network.
@@ 4.1 Dataset
ShowTell |
0.055 |
0.019 |
0.007 |
0.003 |
0.038 |
0.081 |
0.004 |
ShowAttTell |
0.106 |
0.015 |
0.000 |
0.000 |
0.026 |
0.026 |
0.049 |
1NN-Im |
0.071 |
0.020 |
0.007 |
0.004 |
0.032 |
0.069 |
0.059 |
Seq2seq |
0.050 |
0.012 |
0.003 |
0.000 |
0.024 |
0.065 |
0.034 |
1NN-Usr |
0.063 |
0.014 |
0.002 |
0.000 |
0.028 |
0.059 |
0.025 |
1NN-UsrIm |
0.106 |
0.032 |
0.011 |
0.005 |
0.045 |
0.104 |
0.084 |
CSMN-P5 |
0.171 |
0.068 |
0.029 |
0.013 |
0.064 |
0.177 |
0.214 |
CSMN-P5-Mul |
0.145 |
0.049 |
0.022 |
0.009 |
0.049 |
0.145 |
0.143 |
Ours |
0.161 |
0.084 | 0.050 | 0.032 | 0.097 |
0.160 |
0.343 |
@@
ShowTell |
0.027 |
0.003 |
0.000 |
0.000 |
0.024 |
0.043 |
0.003 |
ShowAttTell |
0.088 |
0.010 |
0.001 |
0.000 |
0.034 |
0.116 |
0.076 |
1NN-Im |
0.033 |
0.006 |
0.001 |
0.000 |
0.020 |
0.043 |
0.063 |
Seq2seq |
0.076 |
0.010 |
0.000 |
0.000 |
0.034 |
0.066 |
0.069 |
1NN-Usr |
0.032 |
0.003 |
0.001 |
0.000 |
0.016 |
0.051 |
0.028 |
1NN-UsrIm |
0.039 |
0.005 |
0.001 |
0.000 |
0.021 |
0.050 |
0.050 |
CSMN-P5 |
0.106 |
0.034 |
0.012 |
0.004 |
0.033 |
0.099 |
0.064 |
CSMN-P5-Mul |
0.116 |
0.036 |
0.010 |
0.003 |
0.036 |
0.111 |
0.060 |
Ours |
0.100 |
0.042 | 0.018 | 0.008 | 0.064 |
0.097 |
0.170 |
@@ 4.2 Implementation Details
w/o Context |
0.074 |
0.030 |
0.013 |
0.006 |
0.052 |
0.075 |
0.121 |
w/o Mapping |
0.083 |
0.034 |
0.015 |
0.007 |
0.060 |
0.088 |
0.154 |
w/o Fusion |
0.085 |
0.036 |
0.016 |
0.007 |
0.061 |
0.088 |
0.161 |
w/o Query |
0.083 |
0.035 |
0.015 |
0.006 |
0.061 |
0.089 |
0.152 |
Full Model |
0.100 | 0.042 | 0.018 | 0.008 | 0.064 | 0.097 | 0.170 |

### v2

@@ 2.1 Image Captioning
Image captioning serves as a pivotal link between Computer Vision (CV) and Natural Language Processing (NLP), necessitating the generation of contextually and grammatically appropriate textual descriptions for images. Notable progress has been achieved in this field with the advent of deep learning techniques. Earlier approaches utilized convolutional neural network [7] to extract visual features from images and recurrent neural network [14] to generate captions [13, 28]. Subsequent studies introduced attention mechanisms, leading to improved caption quality [1, 31, 32]. In addition, there has been an emergence of large-scale vision and language pre-trained models that can be applied to image captioning [30, 34]. However, these models often require significant computational resources and training efforts. To address this issue, Mokady et al. [15] proposed a strategy of keeping the pre-trained weights fixed and focusing on training specific network components to bridge the gap between visual and textual representations. Despite the achievements of these efforts, they are limited to generating neutral captions, which may not fully cater to the diverse needs of users in the era characterized by the rapid emergence of social media.
@@ 2.2 Personalized Vision and Language Research
Driven by the rapid proliferation of social media and the growing demand for personalized content, numerous scholars have conducted research with a primary focus on customizing the visual and linguistic domains. Shuster et al. [23] defined 215 potential personality traits to guide caption generation. However, in real-world scenarios, users’ styles may not be adequately captured by these traits. Park et al. [17] constructed datasets specifically for personalized image captioning and employed the memory network to store the users’ active vocabularies, enabling the generation of personalized captions. Long et al. [11] proposed a model for personalized image captioning that can be applied across different domains, thereby expanding the range of potential application scenarios. Considering users’ long-term and short-term stylistic preferences, Zhang et al. [35] proposed an approach that combines these temporal dimensions. However, the implementation of this approach necessitates the collection of more detailed data. To generate captions that align with user preferences, Wang et al. [29] utilized user-provided tags as a foundational element in their methodology. Zeng et al. [33] leveraged user profiles such as age and gender to generate social media comments. Our approach eliminates the need for supplementary information, such as user profiles or posting timestamps, and only requires training a small network.
@@ 4.1 Dataset
| ShowTell | 0.055 | 0.019 | 0.007 | 0.003 | 0.038 | 0.081 | 0.004 |
| ShowAttTell | 0.106 | 0.015 | 0.000 | 0.000 | 0.026 | 0.026 | 0.049 |
| 1NN-Im | 0.071 | 0.020 | 0.007 | 0.004 | 0.032 | 0.069 | 0.059 |
| Seq2seq | 0.050 | 0.012 | 0.003 | 0.000 | 0.024 | 0.065 | 0.034 |
| 1NN-Usr | 0.063 | 0.014 | 0.002 | 0.000 | 0.028 | 0.059 | 0.025 |
| 1NN-UsrIm | 0.106 | 0.032 | 0.011 | 0.005 | 0.045 | 0.104 | 0.084 |
| CSMN-P5 | 0.171 | 0.068 | 0.029 | 0.013 | 0.064 | 0.177 | 0.214 |
| CSMN-P5-Mul | 0.145 | 0.049 | 0.022 | 0.009 | 0.049 | 0.145 | 0.143 |
| Ours | 0.161 | 0.084 | 0.050 | 0.032 | 0.097 | 0.160 | 0.343 |
@@
| ShowTell | 0.027 | 0.003 | 0.000 | 0.000 | 0.024 | 0.043 | 0.003 |
| ShowAttTell | 0.088 | 0.010 | 0.001 | 0.000 | 0.034 | 0.116 | 0.076 |
| 1NN-Im | 0.033 | 0.006 | 0.001 | 0.000 | 0.020 | 0.043 | 0.063 |
| Seq2seq | 0.076 | 0.010 | 0.000 | 0.000 | 0.034 | 0.066 | 0.069 |
| 1NN-Usr | 0.032 | 0.003 | 0.001 | 0.000 | 0.016 | 0.051 | 0.028 |
| 1NN-UsrIm | 0.039 | 0.005 | 0.001 | 0.000 | 0.021 | 0.050 | 0.050 |
| CSMN-P5 | 0.106 | 0.034 | 0.012 | 0.004 | 0.033 | 0.099 | 0.064 |
| CSMN-P5-Mul | 0.116 | 0.036 | 0.010 | 0.003 | 0.036 | 0.111 | 0.060 |
| Ours | 0.100 | 0.042 | 0.018 | 0.008 | 0.064 | 0.097 | 0.170 |
@@ 4.2 Implementation Details
| w/o Context | 0.074 | 0.030 | 0.013 | 0.006 | 0.052 | 0.075 | 0.121 |
| w/o Mapping | 0.083 | 0.034 | 0.015 | 0.007 | 0.060 | 0.088 | 0.154 |
| w/o Fusion | 0.085 | 0.036 | 0.016 | 0.007 | 0.061 | 0.088 | 0.161 |
| w/o Query | 0.083 | 0.035 | 0.015 | 0.006 | 0.061 | 0.089 | 0.152 |
| Full Model | 0.100 | 0.042 | 0.018 | 0.008 | 0.064 | 0.097 | 0.170 |
@@ 5 Conclusions
The Gaussian Mixture Model (GMM) is a traditional machine learning model used for tasks such as probabilistic modeling and clustering. Its network structure is relatively straightforward, typically consisting of multiple Gaussian distributions representing different data clusters or categories. During training, the GMM model estimates the parameters (mean and variance) of each Gaussian distribution and the probabilities of data points belonging to each cluster. In image segmentation tasks, the GMM model can be employed to classify image pixels into various categories, such as background and low-grade gliomas. The GMM model offers advantages such as computational efficiency and interpretability, but its effectiveness in modeling complex data with high-dimensional features is limited.
To further elaborate on the GMM model, let’s consider its application in the context of the LGG-MRI-Segmentation dataset. In this case, the GMM is used to model the distribution of pixel intensities in the MRI images. Each Gaussian component in the GMM represents a distinct tissue class or region, such as tumor or normal brain tissue.
The probability density function of the GMM is defined as a weighted sum of K Gaussian densities:
where is the feature vector of a pixel, is the mixing coefficient (weight) of the k-th Gaussian component, and is the Gaussian density function with mean and covariance matrix .
The parameters of the GMM are estimated using the Expectation-Maximization (EM) algorithm, which iteratively alternates between two steps:
E-step: Calculate the posterior probabilities (responsibilities) of each pixel belonging to each Gaussian component based on the current model parameters. M-step: Update the model parameters (means, covariances, and mixing coefficients) to maximize the likelihood of the data given the current responsibilities. Once the GMM is trained, it can be used to segment new MRI images by assigning each pixel to the Gaussian component with the highest posterior probability.
While the GMM is a simple and interpretable model, it has some limitations when applied to complex datasets like LGG-MRI-Segmentation. The assumption of Gaussian distributions for each tissue class may not always hold true, especially in the presence of artifacts, noise, or heterogeneous tumor appearances. Moreover, the GMM does not explicitly model the spatial dependencies between neighboring pixels, which can lead to fragmented or noisy segmentations.
To address these limitations, various extensions and improvements to the GMM have been proposed, such as incorporating spatial priors, using non-Gaussian distributions, or combining the GMM with other models like Markov Random Fields. However, in recent years, deep learning-based approaches have largely outperformed traditional methods like GMM for medical image segmentation tasks, thanks to their ability to learn complex, hierarchical features from large datasets.